# NOCTIS: Novel Object Cyclic Threshold based Instance Segmentation

### v2

@@ 4 Experiments
In this section we first present our experimental setup (Section 4.1), followed by a comparison of our method with the SOTA ones, across the seven core datasets of the BOP 2023 challenge (Section 4.2). Moreover, we perform ablation studies regarding the score components’ choices in Section 4.3. Finally, in the last Section 4.4, we discuss some limitations of NOCTIS.
@@ 4.1 Experimental setup
We evaluate our method on the seven core datasets of the BOP 2023 challenge: LineMod Occlusion (LM-O) (Brachmann et al., 2014); YCB-Video (YCB-V) (Xiang et al., 2018); TUD-L (Hodan et al., 2018); IC-BIN (Doumanoglou et al., 2016); ITODD (Drost et al., 2017); HomebrewedDB (HB) (Kaskman et al., 2019) and T-LESS (Hodan et al., 2017). Overall those datasets contain $132$ household and industrial objects, being textured or untextured, and are symmetric or asymmetric; moreover, they are shown in multiple cluttered scenes with varying occlusion and lighting conditions.
@@
To generate the proposals, we use GSAM 2, with an input text prompt “objects”, comprised of the Grounding-DINO model with checkpoint “Swin-B” and SAM 2 with checkpoint “sam2.1-L”. The corresponding regions of interest (ROIs) are resized to $224\times 224$, while using padding to keep the original size ratios. We use the default “ViT-L” model/checkpoint of DINOv2 (Oquab et al., 2024), for better comparability with previous works Nguyen et al. (2023); Lin et al. (2024); Lu et al. (2025), to extract the visual descriptors as $1024$-dimensional feature vectors ($N_{\mathit{cls}}^{dim}=N_{\mathit{patch}}^{dim}=1024$), where each patch token on the $16\times 16$ grid represents $14\times 14$ pixels.
We use the “PBR-BlenderProc4BOP” pipeline with the same $42$ predefined viewpoints, as described in CNOS (Nguyen et al., 2023, Sections 3.1 and 4.1), to select the templates representing every dataset object. See also the Appendix Section A.1, for a quick comparison between different types of template renderers.
The main code is implemented in Python 3.8 using Numpy (Harris et al., 2020) and PyTorchPaszke et al. (2019) (Version 2.2.1 cuda). To ensure reproducibility, the seed values of all the (pseudo-) random number generators are set to $2025$. The tests were performed on a single Nvidia RTX 4070 12GB graphics card and the average measured time per run, with one run using the same configuration on all the seven datasets, was approximately $90$ minutes or $0.990$ seconds per image.
@@ 4.2 Comparison with the state-of-the-art
We compare our method with the best available results from the leaderboard [^2] https://bop.felk.cvut.cz/leaderboards/segmentation-unseen-bop23/bop-classic-core/; Accessed: 2025-07-28 of the BOP challenges, comprising of the top-$3$ paper-supported methods: CNOS, SAM-6D and NIDS-Net; and the overall top-$3$ ones: “anonymity”, LDSeg and MUSE (November 2024 and July 2025 version); which do not have a paper or code publicly available. CNOS uses proposals from SAM or FastSAM and only the semantic score 3.3 for matching. SAM-6D uses the same proposals and semantic score as CNOS; additionally, it uses an appearance-based and geometric matching score of the single template with the highest semantic score; the latter score utilizes depth information to consider the shapes and sizes of instances during matching. NIDS-Net uses proposals from GSAM and the similarity between the weight adapter refined Foreground Feature Averaging (Kotar et al., 2023) embeddings together with SAM-6D’s appearance score. For the methods “anonymity”, LDSeg and MUSE, no further information nor clear details are available.
In the Table 1 we show the results for NOCTIS and the other methods on all seven datasets and the overall average. We surpass the best established method NIDS-Net by a significant margin of absolute $3.4\%$ mean AP and we are on par with the best one (MUSE, July 2025 version); overcoming the second best undisclosed one by $0.8\%$. To further highlight that our pipeline performs better (in terms of mean AP score) than previous works like NIDS-Net, and its results do not only stem out of a better foundation model; GSAM 2 was replaced with the older GSAM. The results in the last line of Table 1 show that, even when one makes this change, NOCTIS still achieves results comparable to the SOTA by showing an overall mean AP score of $0.513$. Notably, our methodology outperforms the ones that are using depth data as well (see Section 4.4).
@@
In Figure 3 we show some qualitative segmentation results of our method vs. the publicly available ones, where errors in the masks and/or classifications of the proposals are indicated by red arrows. One can clearly see that all the methods have their own strengths and weaknesses. CNOS and SAM-6D, for example, as they are using SAM/FastSAM as a proposal generator, have problems in differentiating between the objects and some of their parts. While NIDS-Net, due to its internal usage of GSAM, does not suffer from the previously mentioned problem, it still produces misclassifications in the form of labeling scene objects wrongly or by producing oversized bounding boxes around correctly identified objects, leading to multiple detections. NOCTIS, on the other hand, suffers less from said problems, but it is still prone to misclassification, like the other methods, when the objects are too similar looking or too close to each other; see columns $4$ (left clamp) and $6$ (bottom electric boxes and right adapters) for reference. Indeed, in the YCB-V dataset, due to the high similarity between the different sized clamps, one of them is easily confused with the other; the same is true for the electric boxes in T-LESS. Besides, the two stacked adapters are confused as one object and correspondingly misclassified. Overall, NOCTIS still displays fewer errors, on average, when compared to the other methods.
@@ 4.3 Ablation studies
Table 2: Ablation studies on the influence of the components on the mean AP metric. In the $\mathit{s}^{\mathit{appe}}$ column, the values represent the corresponding value of $w_{\mathit{appe}}$, as shown in equation equation 4.
@@
In lines 161-6 of Table 2, we show the influence of the different score components on the mean AP metric to justify our architecture. Line 0 shows the result attained by the complete NOCTIS model as shown in Table 1. The combined use of semantic and appearance scores leads to a better mean AP score than just using them alone, as shown in lines 131-3. Line $4$ then shows that the addition of the proposal confidence to the semantic score also increases the performances significantly. Furthermore, in a pure incremental fashion, one can see that adding the confidence to the pipeline in line $3$, resulting in line $5$, and on top of that adding the CT filter with a value of $\delta_{\mathit{CT}}=5$ yields better results, see lines $5$ and $6$. As one can easily notice, the addition of the score components causes some clear performance gains; however, one cannot justify the inclusion of certain score components in our pipeline by the sheer increase in mean absolute AP score they provide on their own, as one is bound to meet diminishing returns at some point. Intuitively, getting a better mean AP score over all datasets is a daunting task, as some changes in the pipeline might result in better gains over some of them, but decreases over others. As a side note, line $1$, when compared to the standard CNOS (see Table 1), emphasizes the importance of proposal mask quality; thus validating our proposal generator of choice, GSAM 2, against the previously used ones, i.e. SAM and FastSAM.
Further ablation studies regarding different CT values; the effect of $w_{\mathit{appe}}$ on the mean AP score; and the usage of different renders can be found in the Appendix Section A.1.
4.4 Limitations
As seen in column $4$ from Figure 3, our method does not perform at its highest when the objects are similar looking but different sized, e.g. all the clamps from YCB-V; or they are untextured (see column $6$), e.g. the industrial models from ITODD. While these issues might be solved by using depth data, it does not seem that easy, as e.g. SAM-6D is still not able to solve this issue.
Further discussion and limitations regarding the memory usage and runtime are present in the Appendix Section A.2.
5 Conclusion
In this paper we presented NOCTIS, a new framework for zero-shot novel object instance segmentation; which leverages the foundation models Grounded-SAM 2 for object proposal generation and DINOv2 for visual descriptor based matching scores. The novelties introduced in Section 3 have proven to be largely effective; indeed, NOCTIS was able to perform better in terms of mean AP than all the other methodologies (barring MUSE July 2025 version), disclosed or not, on the seven core datasets of the BOP 2023 benchmark. This shows that it is not necessary to have overly complicated scores to achieve good performances. We hope that our work can be used as a new standard baseline to improve upon, especially regarding the formulation of a better scoring rule.
Acknowledgment
Appendix A Appendix
A.1 Further ablation studies
@@
Table 3 demonstrates the effects of using different template sources on the mean AP metric. Line $1$ refers to Pyrender as a lightweight/fast renderer (similar to CNOS) and line $2$ to BlenderProc for more photo realistic renders; which was utilized by GigaPose (Nguyen et al., 2024) for their “onboarding” stage. Both render the object floating in an empty (black) space.
Our default “PBR-BlenderProc4BOP” pipeline gave the best results; while it also uses BlenderProc for rendering, it renders objects in random cluttered scenes to make the reference crops even more realistic. Indeed, the crops obtained this way appear more realistic because of the cross-effects of other objects being in the scene (e.g. slight partial occlusions, other objects’ shadows, changes in the ambient light perspective), compared to standalone objects that produce too clean/artificial templates that seem to have a negative influence on the matching.
For example, if one were to consider the same object in two different scenes, one without any other object and the other with an object that is slightly occluding a part of it. The resulting masks would be different, as the one in the “void” would result in a representation of the object that shows all of its parts; the same cannot be said about the other case, as some of its parts would not be present in its mask. As it turns out, these occluded objects can better represent the ones in the cluttered scenes of the query/test images of the BOP challenge.
Table 3: Ablation study on the influence of the adopted template creation technique.
@@
As it can easily be noticed, Table  4 only shows results for fixed values of the CT used across all objects of all datasets; thus one might wonder what would be the effects of using values adapted to each object, e.g. based on their texture complexity or viewpoint variation, rather than having only a static one. While it might look beneficial to adapt the CT value on object-specific characteristics, NOCTIS is a zero‑shot pipeline requiring no further tuning, aligning with BOP’s challenge goals of methods that work out‑of‑the‑box on novel objects. Introducing adaptive thresholds (partly) undermines this zero‑tuning philosophy and adds dataset‑specific hyperparameters that would require further tuning, probably leading into loss of generalization on new object instances. Moreover, the adjusted CT value of an object would most likely be affected by cross influences of the other objects, making NOCTIS incapable of adding new objects on-the-fly, as re-adjusting these values would always be needed. Eventually, it is unclear, from the get-go, on how many different scenarios one would have to test these new varying CT values internally, considering also the required runtime and memory for such an optimization, to obtain a reliable, general and easily scalable model.
@@
Table 5 illustrates how different $w_{\mathit{appe}}$ values affect the mean AP score on two different configurations of our pipeline; where the first one uses only the semantic and appearance score, while the second one represents the full/best pipeline. While the choice of including or not the appearance score is influential (see Section 4.3), changing its weight as shown by the above results does not provide significant gains/losses over the final performances.
@@
A.2 Additional discussion and limitations
Memory usage and runtime
The memory usage for determining object matching scores, for a query image, has two bottlenecks; each of them requires calculating a large matrix at some point. A $N_{{\bm{I}}}^{\mathit{prop}}\times N^{O}\times N^{T}$ matrix for the semantic score before the aggregation, like in CNOS; and a “big” $N_{{\bm{I}}}^{\mathit{prop}}\times N^{O}\times N^{T}\times N^{\mathit{patch}}\times N^{\mathit{patch}}$ internal floating-point matrix for the appearance score. As a side note, one could think that the size of the latter matrix could be the reason why SAM-6D, originally, computed the appearance score only for one template. Cyclic filtering only temporarily adds ca. seven $N_{{\bm{I}}}^{\mathit{prop}}\times N^{O}\times N^{T}\times N^{\mathit{patch}}$ matrices ($5\times\mathit{integer}+1\times\mathit{float}+1\times\mathit{bool}$) to the memory, which is ca. $36\times$ less memory compared to the big one and so negligible.
To reduce the CUDA memory requirements, we make use of mini batches on the $N_{P}$ and $N_{O}$ dimensions; however, this comes with an increase in computation time. Furthermore, until the templates remain the same between different runs; their visual descriptors stay the same after the onboarding stage (see Section 3.1), since DINOv2 is deterministic, they can be stored and cheaply reloaded later.
Other BOP datasets
Our method is only evaluated on the seven core BOP 2023 datasets; this is done due to, on one hand, a lack of other results’ data for the BOP classic Extra datasets (LM (Hinterstoisser et al., 2013); HOPEv1 (Tyree et al., 2022); RU-APC (Rennie et al., 2016); IC-MI (Tejani et al., 2014) and TYO-L (Hodan et al., 2018)); on the other hand, because the BOP 2024 (Nguyen et al., 2025) and 2025 (Tomas Hodan & Fourmy, 2025) challenges are targeting only a detection task. But, as mentioned in Section 4.1, the chosen datasets provide a wide range of different scenes; therefore, their evaluation should still be reliable.
A.3 Additional comments
@@
Zero-shot generalization has been one of the main driving forces behind NOCTIS’ implementation. Moreover, this is an important prerequisite for the ”Model-based 2D segmentation of unseen objects” task of the BOP challenge to have. Indeed, NOCTIS just needs some template views (see Section3.1) for representing any object from any kind of image source (renderer, video frames, hand-made camera images, etc). For example, one could just take their camera to shoot some photos of an (rigid) object from multiple viewpoints, mask out the object (somehow) and feed the masked photos to NOCTIS as template views; afterwards the object can be detected. Thus, both the CAD/3D object models and the fixed viewpoints are not really necessary for our pipeline; as the models are only needed when rendering the object as a template source, while the viewpoints ensure a good overview of the object. Correspondingly, they are relevant only in the context of the BOP challenge for the evaluation of our pipeline according to the provided benchmarks; which means applying NOCTIS outside of the BOP benchmarks can be easily achieved.
@@
While computational efficiency is usually an important factor for real world applications, it is not an evaluation criterion for the BOP task. Given the fact that different hardware configurations heavily influence the runtime (“normal” graphic cards vs. server ones), the total time per image as stated in the BOP challenge rules is not a major factor regarding the quality of the algorithms proposed. Yet, our pipeline would benefit from better hardware/more CUDA memory, as the “Nvidia RTX 4070 12GB” graphic card used for the experiments is just a standard graphic card compared to what the other participants for the same BOP task were using, like e.g. GeForce RTX 3090 24GB (SAM-6D) or V100 16GB (CNOS).

### v4

@@ 4 Experiments
In this section we first present our experimental setup (Sec. 4.1), followed by a comparison of our method with the SOTA ones, across the seven core datasets of the BOP 2023 challenge (Sec. 4.2). Moreover, we perform ablation studies regarding the score components’ choices in Sec. 4.3. Finally, in the last Sec. 4.4, we discuss some limitations of NOCTIS.
@@ 4.1 Experimental setup
We evaluate our method on the seven core datasets of the BOP 2023 challenge: LineMod Occlusion (LM-O) [Brachmann2014]; T-LESS [hodan2017tlessrgbddataset6d]; TUD-L [hodan2018bopbenchmark6dobject]; IC-BIN [doumanoglou2016recovering6dobjectpose]; ITODD [8265467]; HomebrewedDB (HB) [kaskman2019homebreweddbrgbddataset6d] and YCB-Video (YCB-V) [Xiang-RSS-18]. Overall those datasets contain $132$ household and industrial objects, being textured or untextured, and are symmetric or asymmetric; moreover, they are shown in multiple cluttered scenes with varying occlusion and lighting conditions.
@@
To generate the proposals, we use GSAM 2, with an input text prompt “objects”, comprised of the Grounding-DINO model with checkpoint “Swin-B” and SAM 2 with checkpoint “sam2.1-L”. The corresponding regions of interest (ROIs) are resized to $224\times 224$, while using padding to keep the original size ratios. We use the default “ViT-L” model/checkpoint of DINOv2 [oquab2024dinov2learningrobustvisual], for better comparability with previous works [nguyen2023cnosstrongbaselinecadbased, Lin_2024_CVPR, lu2025adaptingpretrainedvisionmodels], to extract the visual descriptors as $1024$-dimensional feature vectors ($N_{\mathit{cls}}^{dim}=N_{\mathit{patch}}^{dim}=1024$), where each patch token on the $16\times 16$ grid represents $14\times 14$ pixels.
We use the “PBR-BlenderProc4BOP” pipeline with the same $42$ predefined viewpoints, as described in CNOS [nguyen2023cnosstrongbaselinecadbased, Sections 3.1 and 4.1], to select the templates representing every dataset object. See also Sec. 4.3, for a quick comparison between different types of template renderers.
The main code is implemented in Python 3.8 using Numpy [harris2020array] and PyTorch[paszke2019pytorchimperativestylehighperformance] (Version 2.2.1 CUDA 11.8). To ensure reproducibility, the seed values of all the (pseudo-) random number generators are set to $2025$. The tests were performed on a single Nvidia RTX 4070 12GB graphics card and the average measured time per run, with one run using the same configuration on all the seven datasets, was approximately $90$ minutes or $0.990$ seconds per image.
@@ 4.2 Comparison with the state-of-the-art
We compare our method with the best available results from the leaderboard [^2] https://bop.felk.cvut.cz/leaderboards/segmentation-unseen-bop23/bop-classic-core/; Accessed: 2025-05-06 of the BOP challenges, comprising of the top-$3$ paper-supported methods: CNOS, SAM-6D and NIDS-Net; and the overall top-$3$ ones: “anonymity”, LDSeg and MUSE (November 2024 version); which do not have a paper or code publicly available. CNOS uses proposals from SAM or FastSAM and only the semantic score Sec. 3.3 for matching. SAM-6D uses the same proposals and semantic score as CNOS; additionally, it uses an appearance-based and geometric matching score of the single template with the highest semantic score; the latter score utilizes depth information to consider the shapes and sizes of instances during matching. NIDS-Net uses proposals from GSAM and the similarity between the weight adapter refined Foreground Feature Averaging [NEURIPS2023_803c6ab3] embeddings together with SAM-6D’s appearance score. For the methods “anonymity”, LDSeg and MUSE, no further information nor clear details are available.
In the Tab. 1 we show the results for NOCTIS and the other methods on all seven datasets and the overall average. We surpass the best established method NIDS-Net by a significant margin of absolute $3.4\%$ mean AP and the best undisclosed one by $0.8\%$. To further highlight that our pipeline performs better (in terms of mean AP score) than previous works like NIDS-Net, and that its results do not only stem out of a better foundation model; GSAM 2 was replaced with the older GSAM. The results in the last line of Tab. 1 show that, even when one makes this change, NOCTIS still achieves results comparable to the SOTA by showing an overall mean AP score of $0.513$. Notably, our methodology outperforms the ones that are using depth data as well (see Sec. 4.4).
@@
In Fig. 3 we show some qualitative segmentation results of our method vs. the publicly available ones, where errors in the masks and/or classifications of the proposals are indicated by red arrows. One can clearly see that all the methods have their own strengths and weaknesses. CNOS and SAM-6D, for example, as they are using SAM/FastSAM as a proposal generator, have problems in differentiating between the objects and some of their parts. While NIDS-Net, due to its internal usage of GSAM, does not suffer from the previously mentioned problem, it still produces misclassifications in the form of labeling scene objects wrongly or by producing oversized bounding boxes around correctly identified objects, leading to multiple detections. NOCTIS, on the other hand, suffers less from said problems, but it is still prone to misclassification, like the other methods, when the objects are too similar looking or too close to each other; see columns $4$ (left clamp) and $6$ (bottom electric boxes and right adapters) for reference. Indeed, in the YCB-V dataset, due to the high similarity between the different sized clamps, one of them is easily confused with the other; the same is true for the electric boxes in T-LESS. Besides, the two stacked adapters are confused as one object and correspondingly misclassified. Overall, NOCTIS still displays fewer errors, on average, when compared to the other methods.
@@ 4.3 Ablation studies
Table 2: Ablation studies on the influence of the components on the mean AP metric. In the $\mathit{s}^{\mathit{appe}}$ column, the values represent the corresponding value of $w_{\mathit{appe}}$, as shown in Eq. 4.
@@
In lines 161-6 of Tab. 2, we show the influence of the different score components on the mean AP metric to justify our architecture. Line 0 shows the result attained by the complete NOCTIS model as shown in Tab. 1. The combined use of semantic and appearance scores leads to a better mean AP score than just using them alone, as shown in lines 131-3. Line $4$ then shows that the addition of the proposal confidence to the semantic score also increases the performances significantly. Furthermore, in a pure incremental fashion, one can see that adding the confidence to the pipeline in line $3$, resulting in line $5$, and on top of that adding the CT filter with a value of $\delta_{\mathit{CT}}=5$ yields better results, see lines $5$ and $6$. As one can easily notice, the addition of the score components causes some clear performance gains; however, one cannot justify the inclusion of certain score components in our pipeline by the sheer increase in mean absolute AP score they provide on their own, as one is bound to meet diminishing returns at some point. Intuitively, getting a better mean AP score over all datasets is a daunting task, as some changes in the pipeline might result in better gains over some of them, but decreases over others. As a side note, line $1$, when compared to the standard CNOS (see Tab. 1), emphasizes the importance of proposal mask quality; thus validating our proposal generator of choice, GSAM 2, against the previously used ones, i.e. SAM and FastSAM.
@@
Table 3 demonstrates the effects of using different template sources on the mean AP metric. Line $1$ refers to Pyrender as a lightweight/fast renderer (similar to CNOS) and line $2$ to BlenderProc for more photo realistic renders; which was utilized by GigaPose [nguyen2024gigaposefastrobustnovel] for their “onboarding” stage. Both render the object floating in an empty (black) space.
Our default “PBR-BlenderProc4BOP” pipeline gave the best results; while it also uses BlenderProc for rendering, it renders objects in random cluttered scenes to make the reference crops even more realistic. Indeed, if one were to consider the same object in two different scenes, one without any other object and the other with an object that is slightly occluding a part of it, the resulting masks would be different. The one in the “void” would result in a too clean/artificial representation of the object that shows all of its parts, while the other would not have some of its parts in its mask. As it turns out, these occluded objects can better represent the ones in the cluttered scenes of the query/test images of the BOP challenge. Further ablation studies regarding different CT values and the effect of $w_{\mathit{appe}}$ on the mean AP score can be found in the Appendix Sec. 6.1.
Table 3: Ablation study on the influence of the adopted template creation technique on the mean AP metric.
@@
4.4 Limitations
Objects’ size
As seen in column $4$ from Fig. 3, our method does not perform at its highest when the objects are similar looking but different sized, e.g. all the clamps from YCB-V; or they are untextured (see column $6$), e.g. the industrial models from ITODD. While these issues might be solved by using depth data, it does not seem that easy, as e.g. SAM-6D is still not able to solve this issue.
Memory usage and runtime
The memory usage for determining object matching scores, for a query image, has two bottlenecks; each of them requires calculating a large matrix at some point. A $N_{{\bm{I}}}^{\mathit{prop}}\times N^{O}\times N^{T}$ matrix for the semantic scores before the aggregation, like in CNOS; and a “big” $N_{{\bm{I}}}^{\mathit{prop}}\times N^{O}\times N^{T}\times N^{\mathit{patch}}\times N^{\mathit{patch}}$ internal floating-point matrix for the appearance scores. Cyclic filtering only temporarily adds ca. seven $N_{{\bm{I}}}^{\mathit{prop}}\times N^{O}\times N^{T}\times N^{\mathit{patch}}$ matrices ($5\times\mathit{integer}+1\times\mathit{float}+1\times\mathit{bool}$) to the memory, which is ca. $36\times$ less memory compared to the big one and so negligible.
Other BOP datasets
Our method is only evaluated on the seven core BOP 2023 datasets; this is done due to, on one hand, a lack of other results’ data for the BOP classic Extra datasets (LM [10.1007/978-3-642-37331-2_42]; HOPEv1 [tyree20226dofposeestimationhousehold]; RU-APC [rennie2016datasetimprovedrgbdbasedobject]; IC-MI [10.1007/978-3-319-10599-4_30] and TYO-L [hodan2018bopbenchmark6dobject]); on the other hand, because the BOP 2024 [nguyen2025bopchallenge2024modelbased] and 2025 [bop2025challenges] challenges are targeting only a detection task. But, as mentioned in Section 4.1, the chosen datasets provide a wide range of different scenes; therefore, their evaluation should still be reliable.
5 Conclusion
In this paper we presented NOCTIS, a new framework for zero-shot novel object instance segmentation; which leverages the foundation models Grounded-SAM 2 for object proposal generation and DINOv2 for visual descriptor based matching scores. The novelties introduced in Sec. 3 have proven to be largely effective; indeed, NOCTIS was able to perform better in terms of mean AP than all the other methodologies, disclosed or not, on the seven core datasets of the BOP 2023 benchmark. This shows that it is not necessary to have overly complicated scores to achieve good performances. We hope that our work can be used as a new standard baseline to improve upon, especially regarding the formulation of a better scoring rule.
Acknowledgment
[…]
Supplementary Material
6 Appendix
6.1 Further ablation studies
@@
As it can easily be noticed, Tab. 4 only shows results for fixed values of the CT used across all objects of all datasets; thus one might wonder what would be the effects of using values adapted to each object, e.g. based on their texture complexity or viewpoint variation, rather than having only a static one. While it might look beneficial to adapt the CT value on object-specific characteristics, NOCTIS is a zero‑shot pipeline requiring no further tuning, aligning with BOP’s challenge goals of methods that work out‑of‑the‑box on novel objects. Introducing adaptive thresholds (partly) undermines this zero‑tuning philosophy and adds dataset‑specific hyperparameters that would require further tuning, probably leading into loss of generalization on new object instances. Moreover, the adjusted CT value of an object would most likely be affected by cross influences of the other objects, making NOCTIS incapable of adding new objects on-the-fly, as re-adjusting these values would always be needed. Eventually, it is unclear, from the get-go, on how many different scenarios one would have to test these new varying CT values internally, considering also the required runtime and memory for such an optimization, to obtain a reliable, general and easily scalable model.
@@
Table 5 illustrates how different $w_{\mathit{appe}}$ values affect the mean AP score on two different configurations of our pipeline; where the first one uses only the semantic and appearance score, while the second one represents the full/best pipeline. While the choice of including or not the appearance score is influential (see Sec. 4.3), changing its weight as shown by the above results does not provide significant gains/losses over the final performances.
@@
6.2 Additional comments
@@
Zero-shot generalization has been one of the main driving forces behind NOCTIS’ implementation. Moreover, this is an important prerequisite for the “Model-based 2D segmentation of unseen objects” task of the BOP challenge to have. Indeed, NOCTIS just needs some template views (see Sec. 3.1) for representing any object from any kind of image source (renderer, video frames, hand-made camera images, etc). For example, one could just take their camera to shoot some photos of an (rigid) object from multiple viewpoints, mask out the object (somehow) and feed the masked photos to NOCTIS as template views; afterwards the object can be detected. Thus, both the CAD/3D object models and the fixed viewpoints are not really necessary for our pipeline; as the models are only needed when rendering the object as a template source, while the viewpoints ensure a good overview of the object. Correspondingly, they are relevant only in the context of the BOP challenge for the evaluation of our pipeline according to the provided benchmarks; which means applying NOCTIS outside of the BOP benchmarks can be easily achieved.
@@
While computational efficiency is usually an important factor for real world applications, it is not an evaluation criterion for the BOP task. Given the fact that different hardware configurations heavily influence the runtime (“normal” graphic cards vs. server ones), the total time per image as stated in the BOP challenge rules is not a major factor regarding the quality of the algorithms proposed. Yet, our pipeline would benefit from better hardware/more CUDA memory, as the “Nvidia RTX 4070 12GB” graphic card used for the experiments is just a standard graphic card compared to what the other participants for the same BOP task were using, like e.g. “GeForce RTX 3090 24GB” (SAM-6D) or “V100 16GB” (CNOS).